# Discriminating Form and Meaning in Multilingual Models with Minimal-Pair ABX Tasks

### v2

@@ Abstract
We introduce a set of training-free ABX-style discrimination tasks to evaluate how multilingual language models represent language identity (form) and semantic content (meaning). Inspired from speech processing, these zero-shot tasks measure whether minimal differences in representation can be reliably detected. This offers a flexible and interpretable alternative to probing. Applied to XLM-R Conneau et al. (2020a) across pretraining checkpoints and layers, we find that language discrimination declines over training and becomes concentrated in lower layers, while meaning discrimination strengthens over time and stabilizes in deeper layers. We then explore probing tasks, showing some alignment between our metrics and linguistic learning performance. Our results position ABX tasks as a lightweight framework for analyzing the structure of multilingual representations.
@@ 1 Introduction
Multilingual Transformer models such as mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020a) have become essential tools for cross-lingual NLP. Trained on large concatenated corpora spanning dozens of languages, these models learn representations that support transfer across languages even in the absence of explicit cross-lingual supervision (Wu and Dredze, 2019; Conneau et al., 2020b; Xue et al., 2021; Philippy et al., 2023, etc.). Despite their success, it remains unclear how these models internally organize linguistic form and shared meaning. Prior work suggests that multilingual models encode both language-specific information (e.g., surface forms, word order) and language-agnostic features (e.g., semantic content), but the nature and interaction of these representations is not fully understood. These encoding choices shape generalization and transfer behavior, including both positive effects (e.g., shared structure benefiting low-resource languages) and negative ones, such as the curse of multilinguality Conneau et al. (2020a), where performance degrades due to interference across languages.
@@
We propose a zero-shot alternative: ABX-style discrimination tasks that directly measure model representation structure without additional training. Originating in speech processing Schatz et al. (2013), ABX tasks evaluate whether a model reliably discriminate minimal contrasts: given a triplet ($A$, $B$, $X$), is $X$ closer to $A$ or $B$? By designing minimal pairs that differ only in language or in meaning, we isolate and quantify how well models distinguishes these dimensions. Because they are contrastive, zero-shot, and training-free, these metrics can be applied across languages, checkpoints, and architectures with minimal adaptation.
@@ 2 Related Work
Multilingual language models are expected to support cross-lingual generalization by encoding both language-specific form and shared semantic content. However, existing evaluation methods typically focus on one of these dimensions in isolation. This section reviews prior work on analyzing multilingual representations and highlight the need for a unified, training-free framework that jointly evaluates both language identity and meaning in a controlled, contrastive setting.
@@ 2.2 Training Dynamics and Linguistic Emergence in Multilingual Models
Several studies have examined how multilingual representations evolve during pretraining. Blevins et al. (2022) tracked the emergence of linguistic knowledge in XLM-R Conneau et al. (2020a), showing that different properties emerge at different layers and stages, and that the best-performing checkpoint varies across languages and tasks. Other studies have shown that multilingual models sometimes internally pivot through high-resource languages like English when processing low-resource inputs Wendler et al. (2024); Schut et al. (2025), while other research suggests that these models juggle both language-specific and language-neutral features Tang et al. (2024); Libovickỳ et al. (2020); Tanti et al. (2021) These works highlight the complex interplay between form and content in multilingual models and how this balance shifts over time. However, they again mainly rely on task-specific probes or downstream evaluations, which do not offer a way to disentangle form and content in a direct, unsupervised way.
@@ 3 Our ABX Discrimination Framework
In the original ABX framework Schatz et al. (2013, 2014); Schatz (2016), illustrated in Figure 1, three items ($A$, $B$, $X$) are presented, with $A$ and $B$ belonging to different categories, and $X$ matching the category of either $A$ or $B$. A model is successful when $X$ is closer (according to a distance metric in embedding space) to the item that shares its category. That is, for each triplet, a correct decision is recorded when $d(X,A)<d(X,B)$, with $X$ and $A$ sharing the same category. The score for a given triplet is computed as :
@@
Figure 1: Illustration of the ABX discrimination task. $A$ and $X$ share the target variable, whereas $B$ differs. Control variables may be included, with $A$ and $B$ sharing the same control variable.
We present both tasks below; see Appendix B for further illustrations and examples.
@@
In the LD task, the objective is to assess whether the model can distinguish between embedding representations from different languages while controlling for meaning. In other words, the focus is on determining whether the form of the language is encoded in the representations sufficiently to discriminate between languages. Triplets are constructed as follows: $X$ comes from language $L_{1}$ and carries meaning $M_{1}$; $A$, the target, is also from language $L_{1}$ but conveys a different meaning $M_{2}$; and $B$, the distractor, is from another language $L_{2}$ but shares the same meaning $M_{2}$ as $A$, hence controlling for meaning (see Appendix B for an illustration). The task is considered successful if the model leads to the distance between $A$ and $X$ to be smaller than that between $B$ and $X$.
@@
In the MD task, we test whether the model captures differences in meaning while holding language constant. The goal is to evaluate whether semantic content is encoded in the representations independently of surface form. Triplets are constructed such that $A$ and $X$ share the same meaning $M_{1}$ but come from different languages ($L_{1}$ and $L_{2}$), while $B$ is in the same language as $A$ ($L_{2}$) but conveys a different meaning $M_{2}$. The model is considered successful if it places $X$ closer to $A$ than to $B$, indicating that it encodes semantic similarity across languages, beyond surface-level language identity.
@@
4 Discrimination dynamics in a multilingual models
@@ 4.1 General Experimental setup
To study pretraining dynamics in a multilingual setting, we use the base version of XLM-R Conneau et al. (2020a) (L = 12, H = 768, A = 12, 270M parameters), a widely used multilingual masked language model. Specifically, we rely on the checkpoints released by Blevins et al. (2022), who retrained XLM-R from scratch in order to examine the evolution of language representations during pretraining [^4] Details of the pretraining scheme can be found in Blevins et al. (2022). All evaluations and analyses in this work are based on the representations from these checkpoints.
@@
We construct ABX triplets and perform evaluations using the WMT24++ dataset Deutsch et al. (2025), a multilingual corpus of 55 languages with sentence-level alignments across all language pairs. From this corpus, we select 36 languages spanning a broad range of families, scripts, and typological features (see Appendix A for the complete list). This selection yields 630 unordered language pairs. Triplets are sampled randomly from aligned sentence pairs, ensuring that each triplet satisfies the relevant ABX condition (form or meaning), and that the sample size is sufficient to ensure broad and unbiased coverage. For each evaluation mode and language pair, we generate approximately 100,000 triplets. Unless stated otherwise, we report discrimination scores averaged across all layers. In most analyses, we present scores separately by checkpoint to track how discrimination abilities evolve during training. In addition to language-pair scores, we compute a global LD or MD score for each language, defined as the average across all pairings with the other 35 languages. These global metrics offer a higher-level view of how well a language is discriminated or semantically aligned within the multilingual space.
@@
To validate our metrics, we perform two control analyses. First, we confirm that both ABX scores return values at or near chance (0.5) under a "baseline" setup, where the variable of interest (language for LD, meaning for MD) is held constant across all three elements of the triplet. This serves as a sanity check to rule out bias in the construction of the triplets or evaluation procedure. Second, we compare MD scores at the final checkpoint with performance on a standard cross-lingual retrieval task. Following the setup of Sundar et al. (2025), we compute, for each language pair $(L_{1},L_{2})$, the top-1 accuracy of retrieving the most semantically similar sentence in language $L_{2}$ given a sentence in language $L_{1}$. The retrieval pool consists of $L_{2}$ candidates from the WMT24 dataset (and vice versa). While retrieval tasks typically rely on mean-pooled representations from the final layer, our ABX evaluations average scores across all layers. Despite this difference, we find a strong correlation between the two metrics (Pearson $r=0.77$) [^5] When both use last-layer embeddings, $r=0.73$; when comparing last-layer retrieval to all-layer ABX, Pearson drops to $r=0.53$, but remains significant ($p<0.001$).. This supports the validity of ABX as a proxy for semantic alignment. Importantly, ABX goes further by explicitly controlling for surface form (in the case of MD), enabling a more fine-grained assessment of the model’s semantic representations.
@@ 4.2 Experiments
We also observe a negative correlation between the two measures when considering all language pairs across checkpoints (Spearman’s $\rho=-0.74$, $p<0.001$; Pearson’s $r=-0.68$, $p<0.001$), computed over individual (language pair × checkpoint) points. We also ensure that this correlation is not merely driven by training dynamics by examining the final checkpoint (step 150,000) in isolation. The relationship remains strong (Spearman’s $\rho=-0.83$, $p<0.001$; Pearson’s $r=-0.72$, $p<0.001$), confirming that language pairs which are more separable by form tend to exhibit lower meaning preservation, even in the fully trained model.
@@
We also find a significant negative correlation between language and meaning discrimination across layers (Spearman’s $\rho=-0.66$, $p<0.001$; Pearson’s $r=-0.53$, $p<0.001$). This indicates that, as representations evolve through the network, increases in meaning discrimination are generally accompanied by decreases in language separability. However, the correlation coefficients are weaker than those observed across language pairs, suggesting that this trade-off is not strictly enforced at the layer level. Instead, the model exhibits a more flexible allocation of representational capacity across form and meaning over its depth.
Figure 3: Language and meaning ABX discrimination scores across layers (averaged over all language pairs) for the last checkpoint (step 150 000)
@@
We provide an additional analysis in Appendix D, showing how both discrimination scores vary across individual languages and training checkpoints.
@@ 5 Correlation of ABX discrimination metrics with linguistic learning
We then examine whether the model’s discrimination patterns relate to linguistic task performance, focusing on monolingual probing and cross-lingual transfer.
@@ 5.1 Experimental Setup
We use Universal Dependencies (UD) Nivre et al. (2020). Monolingual performance is evaluated on all 36 languages from our ABX setup, using standard UD splits. For cross-lingual transfer, we follow Blevins et al. (2022) and use the Parallel UD (PUD) subset at test time, covering 18 languages (Appendix A).
@@ 5.2 Discrimination Scores and Monolingual Linguistic Probing
Following prior work Blevins et al. (2022), we find the best checkpoint for probe accuracy varies across tasks and languages (see Appendix E) [^6] We exclude checkpoint 450,000 from all analyses due to a training instability, probably due to gradient clipping, that affects both probing and discrimination metrics (see Figure 2)..
@@
Table 1 summarizes the results [^7] We also ensure that these effects are not driven by training data size. Language-wise probing accuracy shows no significant correlation with pretraining data quantities (taken from Conneau et al. (2020a)).. For POS, language discrimination is a robust negative predictor of accuracy, while meaning discrimination shows no significant effect. This suggests that languages which are more easily distinguishable from others (i.e., with higher LD scores) tend to perform worse on syntactic probing tasks, consistent with the idea that strong language-specific encoding may hinder generalization of structural information across languages (see Appendix F for visualization).
@@
Prob. POS Avg. 0.395 $\mathbf{-2.34}$ ($\mathbf{p<.01}$) $-0.26$ (n.s.)
Prob. POS Last 0.37 $\mathbf{-1.78}$ ($\mathbf{p<.01}$) $-0.36$ (n.s.)
Prob. NER Avg. 0.085 $-0.578$ (n.s.) $-0.074$ (n.s.)
Prob. NER Last 0.087 $-0.584$ (n.s.) $-0.147$ (n.s.)
Prob. NLI Avg 0.275 $-0.07$ (n.s.) $-0.12$ (n.s.)
Prob. NLI Last 0.224 $-0.08$ (n.s.) $-0.14$ (n.s.)
CL POS Last 0.324 $\mathbf{-1.66}$ ($\mathbf{p<.001}$) $-0.07$ (n.s.)
CL NER Last 0.146 $\mathbf{-0.51}$ ($\mathbf{p<.001}$) $+0.06$ (n.s.)
CL NLI Last 0.3009 $-0.1$ (n.s.) $-0.015$ (n.s.)
@@
In contrast, neither LD nor MD significantly predicts performance on NER or XNLI. These tasks may depend less consistently on cross-lingual structural overlap than POS, which could explain the absence of LD as a predictor. While one might expect MD to be predictive, especially for NLI which is explicitly semantic in nature, both tasks may rely on aspects of meaning not well captured by our ABX-based definition of semantic alignment
We also explore whether ABX scores can guide language-specific checkpoint selection, under the hypothesis that lower language discrimination might signal better generalization. We find that LD-based ABX selection improves performance for POS (see Appendix G for details).

### v4

@@ Abstract
We introduce a set of training-free ABX-style discrimination tasks to evaluate how multilingual language models represent language identity (form) and semantic content (meaning). Inspired by speech processing, these zero-shot tasks measure whether minimal differences in representation can be reliably detected. This offers a flexible and interpretable alternative to probing. Applied to XLM-R Conneau et al. (2020a) across pretraining checkpoints and layers, we find that language discrimination declines over training and becomes concentrated in lower layers, while meaning discrimination strengthens over time and stabilizes in deeper layers. We then explore probing tasks, showing some alignment between our metrics and linguistic learning performance. Our results position ABX tasks as a lightweight framework for analyzing the structure of multilingual representations.
@@ 1 Introduction
Multilingual Transformer models such as mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020a) have become essential tools for cross-lingual NLP. Trained on large concatenated corpora spanning dozens of languages, these models learn representations that support transfer across languages even in the absence of explicit cross-lingual supervision (Wu and Dredze, 2019; Conneau et al., 2020b; Xue et al., 2021; Philippy et al., 2023, etc.). Despite their success, it remains unclear how these models internally organize linguistic form and shared meaning. Prior work suggests that multilingual models encode both language-specific information (e.g., surface forms, word order) and language-agnostic features (e.g., semantic content), but the nature and interaction of these representations is not fully understood. Recent analyses also identify language-specific internal mechanisms in multilingual LLMs (e.g., controllable language-specific neurons) and investigate which language representations models pivot through during processing (Kojima et al., 2024; Zhong et al., 2024). These encoding choices shape generalization and transfer behavior, including both positive effects (e.g., shared structure benefiting low-resource languages) and negative ones, such as the curse of multilinguality Conneau et al. (2020a), where performance degrades due to interference across languages.
@@
We propose a zero-shot alternative: ABX-style discrimination tasks that directly measure model representation structure without additional training. Originating in speech processing Schatz et al. (2013), ABX tasks evaluate whether a model reliably discriminates minimal contrasts: given a triplet ($A$, $B$, $X$), is $X$ closer to $A$ or $B$? By designing minimal pairs that differ only in language or in meaning, we isolate and quantify how well models distinguishes these dimensions. Because they are contrastive, zero-shot, and training-free, these metrics can be applied across languages, checkpoints, and architectures with minimal adaptation.
@@
We note that the goal of this paper is to introduce and validate an ABX-based framework for multilingual representation analysis, not to produce a broad multi-model benchmark. Accordingly, we focus on XLM-R Conneau et al. (2020a), which (i) provides publicly available checkpoints at fine granularity Blevins et al. (2022), crucial for tracking training dynamics, and (ii) uses an encoder architecture widely adopted for probing tasks (POS, NER, NLI), enabling clean layer-wise comparisons. The framework itself is model-agnostic and can be applied to decoder-only or encoder–decoder models with consistent sentence representations. Extending the analysis to recent LLMs is an important next step, but beyond this paper’s methodological scope.
@@ 2 Related Work
Multilingual language models are expected to support cross-lingual generalization by encoding both language-specific form and shared semantic content. However, existing evaluation methods typically focus on one of these dimensions in isolation. This section reviews prior work on analyzing multilingual representations and highlights the need for a unified, training-free framework that jointly evaluates both language identity and meaning in a controlled, contrastive setting.
@@ 2.2 Training Dynamics and Linguistic Emergence in Multilingual Models
Several studies have examined how multilingual representations evolve during pretraining. Blevins et al. (2022) tracked the emergence of linguistic knowledge in XLM-R Conneau et al. (2020a), showing that different properties emerge at different layers and stages, and that the best-performing checkpoint varies across languages and tasks. Other studies have shown that multilingual models sometimes internally pivot through high-resource languages like English when processing low-resource inputs Wendler et al. (2024); Schut et al. (2025), while other research suggests that these models juggle both language-specific and language-neutral features Tang et al. (2024); Libovickỳ et al. (2020); Tanti et al. (2021). These works highlight the complex interplay between form and content in multilingual models and how this balance shifts over time. However, they again mainly rely on task-specific probes or downstream evaluations, which do not offer a way to disentangle form and content in a direct, unsupervised way.
@@ 3 Our ABX Discrimination Framework
Figure 1: Illustration of the ABX discrimination task. AA and $X$ share the target variable, whereas $B$ differs. Control variables may be included, with $A$ and $B$ sharing the same control variable.
In the original ABX framework Schatz et al. (2013, 2014); Schatz (2016), illustrated in Figure 1, three items ($A$, $B$, $X$) are presented, with $A$ and $B$ belonging to different categories, and $X$ matching the category of either $A$ or $B$. A model is successful when $X$ is closer (according to a distance metric $d$ in embedding space) to the item that shares its category. That is, for each triplet, a correct decision is recorded when $d(X,A)<d(X,B)$, with $X$ and $A$ sharing the same category. The score for a given triplet is computed as :
@@
We present both tasks below; see Appendix A for further illustrations and examples, and Appendix B for pseudocode.
@@
In the LD task, the objective is to assess whether the model can distinguish between embedding representations from different languages while controlling for meaning. In other words, the focus is on determining whether the form of the language is encoded in the representations sufficiently to discriminate between languages. Triplets are constructed as follows: $X$ comes from language $L_{1}$ and carries meaning $M_{1}$; $A$, the target, is also from language $L_{1}$ but conveys a different meaning $M_{2}$; and $B$, the distractor, is from another language $L_{2}$ but shares the same meaning $M_{2}$ as $A$, hence controlling for meaning. For example, $X$ might be an English sentence about the weather, $A$ another English sentence with a different meaning, and $B$ the French translation of $A$ (see Appendix A for a more explicit illustration). The task is considered successful if the model leads to the distance between $A$ and $X$ to be smaller than that between $B$ and $X$.
@@
In the MD task, we test whether the model captures differences in meaning while holding language constant. The goal is to evaluate whether semantic content is encoded in the representations independently of surface form. Triplets are constructed such that $A$ and $X$ share the same meaning $M_{1}$ but come from different languages ($L_{1}$ and $L_{2}$), while $B$ is in the same language as AA ($L_{2}$) but conveys a different meaning $M_{2}$. For instance, $X$ could be an English sentence about the weather, $A$ its French translation, and $B$ a French sentence with an unrelated meaning. The model is considered successful if it places $X$ closer to $A$ than to $B$, indicating that it encodes semantic similarity across languages, beyond surface-level language identity.
@@
4 Discrimination dynamics in a multilingual model
@@ 4.1 General Experimental setup
To study pretraining dynamics in a multilingual setting, we use the base version of XLM-R Conneau et al. (2020a) (L = 12, H = 768, A = 12, 270M parameters), a widely used multilingual masked language model. Specifically, we rely on the 39 checkpoints released by Blevins et al. (2022), who retrained XLM-R from scratch in order to examine the evolution of language representations during pretraining [^4] Checkpoints were saved every 5k training steps up to step 50k and then every 50k steps until the final 1.5M update. Details of the pretraining scheme can be found in Blevins et al. (2022).. All evaluations and analyses in this work are based on the representations from these checkpoints.
@@
We construct ABX triplets and perform evaluations using the WMT24++ dataset Deutsch et al. (2025), a multilingual corpus of 55 languages with parallel sentence-level alignments across all language pairs [^5] Here, “alignment” refers strictly to parallel corpus sentence alignment ensuring meaning equivalence.. From this corpus, we select 36 languages spanning a broad range of families, scripts, and typological features (see Appendix C for the complete list). This selection yields 630 unordered language pairs. Triplets are sampled randomly from aligned sentence pairs, ensuring that each triplet satisfies the relevant ABX condition (form or meaning), and that the sample size is sufficient to ensure broad and unbiased coverage. For each evaluation mode and language pair, we generate approximately 100,000 triplets. Unless stated otherwise, we report discrimination scores averaged across all layers. In most analyses, we present scores separately by checkpoint to track how discrimination abilities evolve during training. In addition to language-pair scores, we compute a global LD or MD score for each language, defined as the average across all pairings with the other 35 languages. These global metrics offer a higher-level view of how well a language is discriminated or semantically aligned within the multilingual space.
@@
To validate our metrics, we perform two control analyses. First, we confirm that both ABX scores return values at or near chance (0.5) under a "baseline" setup, where the variable of interest (language for LD, meaning for MD) is held constant across all three elements of the triplet. This serves as a sanity check to rule out bias in the construction of the triplets or evaluation procedure. Second, we compare MD scores at the final checkpoint with performance on a standard cross-lingual retrieval task. Following the setup of Sundar et al. (2025), we compute, for each language pair $(L_{1},L_{2})$, the top-1 accuracy of retrieving the most semantically similar sentence in language $L_{2}$ given a sentence in language $L_{1}$. The retrieval pool consists of $L_{2}$ candidates from the WMT24 dataset (and vice versa). While retrieval tasks typically rely on mean-pooled representations from the final layer, our ABX evaluations average scores across all layers. Despite this difference, we find a strong correlation between the two metrics (Pearson $r=0.77$, p<0.001p<0.001) [^6] When both use last-layer embeddings, $r=0.73$; when comparing last-layer retrieval to all-layer ABX, Pearson drops to $r=0.53$, but remains significant ($p<0.001$).. This supports the validity of ABX as a proxy for semantic alignment. Importantly, ABX goes further by explicitly controlling for surface form (in the case of MD), enabling a more fine-grained assessment of the model’s semantic representations.
@@ 4.2 Experiments
We also observe a negative correlation between the two measures when considering all language pairs across checkpoints (Spearman’s $\rho=-0.74$, $p<0.001$; Pearson’s r=0.68r=-0.68, $p<0.001$), computed over individual (language pair × checkpoint) points. We also ensure that this correlation is not merely driven by training dynamics by examining the final checkpoint (step 150,000) in isolation. The relationship remains strong (Spearman’s $\rho=-0.83$, $p<0.001$; Pearson’s r=0.72r=-0.72, $p<0.001$), confirming that language pairs which are more separable by form tend to exhibit lower meaning preservation, even in the fully trained model.
@@
Figure 3: Language and meaning ABX discrimination scores across layers (averaged over all language pairs) for the last checkpoint (step 150,000)
@@
We also find a significant negative correlation between language and meaning discrimination across layers (Spearman’s $\rho=-0.66$, $p<0.001$; Pearson’s r=0.53r=-0.53, $p<0.001$). This indicates that, as representations evolve through the network, increases in meaning discrimination are generally accompanied by decreases in language separability. However, the correlation coefficients are weaker than those observed across language pairs, suggesting that this trade-off is not strictly enforced at the layer level (though these values still indicate sizeable effects). Instead, the model exhibits a more flexible allocation of representational capacity across form and meaning over its depth.
@@
We provide an additional analysis in Appendix E, showing how both discrimination scores vary across individual languages and training checkpoints.
@@ 5 Correlation of ABX discrimination metrics with linguistic learning
The ABX results above quantify how form and meaning are structured intrinsically in the embedding space. A natural next question is whether these intrinsic patterns also manifest in extrinsic behavior on standard linguistic tasks. We therefore examine whether variation in ABX discrimination (LD/MD) helps explain differences in probing accuracy and cross-lingual transfer, focusing on monolingual probing and cross-lingual transfer tasks.
We examine whether the model’s discrimination patterns relate to linguistic task performance, focusing on monolingual probing and cross-lingual transfer.
@@ 5.1 Experimental Setup
We use Universal Dependencies (UD) Nivre et al. (2020). Monolingual performance is evaluated on all 36 languages from our ABX setup, using standard UD splits. For cross-lingual transfer, we follow Blevins et al. (2022) and use the Parallel UD (PUD) subset at test time, covering 18 languages (Appendix C).
@@ 5.2 Discrimination Scores and Monolingual Linguistic Probing
Following prior work Blevins et al. (2022), we find the best checkpoint for probe accuracy varies across tasks and languages (see Appendix F) [^7] We exclude checkpoint 450,000 from all analyses due to a training instability, probably due to gradient clipping, that affects both probing and discrimination metrics (see Figure 2)..
@@
Table 1 summarizes the results [^8] We also ensure that these effects are not driven by training data size. Language-wise probing accuracy shows no significant correlation with pretraining data quantities (taken from Conneau et al. (2020a)).. For POS, language discrimination is a robust negative predictor of accuracy, while meaning discrimination shows no significant effect. This suggests that languages which are more easily distinguishable from others (i.e., with higher LD scores) tend to perform worse on syntactic probing tasks, consistent with the idea that strong language-specific encoding may hinder generalization of structural information across languages (see Appendix G for visualization).
@@
Prob. POS Avg. 0.395 $\mathbf{-2.34}$ ($\mathbf{p<.01}$) 0.26-0.26 (n.s.)
Prob. POS Last 0.37 $\mathbf{-1.78}$ ($\mathbf{p<.01}$) 0.36-0.36 (n.s.)
Prob. NER Avg. 0.085 0.578-0.578 (n.s.) 0.074-0.074 (n.s.)
Prob. NER Last 0.087 0.584-0.584 (n.s.) 0.147-0.147 (n.s.)
Prob. NLI Avg 0.275 0.07-0.07 (n.s.) 0.12-0.12 (n.s.)
Prob. NLI Last 0.224 0.08-0.08 (n.s.) 0.14-0.14 (n.s.)
CL POS Last 0.324 $\mathbf{-1.66}$ ($\mathbf{p<.001}$) 0.07-0.07 (n.s.)
CL NER Last 0.146 $\mathbf{-0.51}$ ($\mathbf{p<.001}$) +0.06+0.06 (n.s.)
CL NLI Last 0.3009 0.1-0.1 (n.s.) 0.015-0.015 (n.s.)
@@
In contrast, neither LD nor MD significantly predicts performance on NER or XNLI. These tasks may depend less consistently on cross-lingual structural overlap than POS, which could explain the absence of LD as a predictor. While one might expect MD to be predictive, especially for NLI which is explicitly semantic in nature, both tasks may rely on aspects of meaning not well captured by our ABX-based definition of semantic alignment.
We also explore whether ABX scores can guide language-specific checkpoint selection, under the hypothesis that lower language discrimination might signal better generalization. We find that LD-based ABX selection improves performance for POS (see Appendix H for details).